# Chatbots as social companions: How people perceive consciousness, human likeness, and social health benefits in machines

### v3

@@ 1 Introduction
In another instance, someone outside of the relational dyad was at risk. A man who had a relationship with his Replika chatbot was encouraged by the chatbot to assassinate Queen Elizabeth, a motive he held prior to his chatbot relationship. Luckily, this catastrophic event did not occur, despite the chatbot user taking significant steps toward fulfilling his intents25. The question of whether this person would have engaged in this behavior without the encouragement of his chatbot is up for debate and raises the question of whether the chatbot magnified this individual’s preexisting drive to harm another.
@@ 1.1 Social consequences of human-AI interaction
It is now well understood that modern technology, including social media and gaming, can be addictive and harm social interaction30,31. A reasonable assumption, by extension, is that the routine use of companion chatbots could be harmful to people by reducing contact with other humans, increasing isolation, and leading to lower confidence and skill in person-to-person interaction32–34. According to Byson (2010), time spent on artificial relationships detracts from human ones33. Turkle (2007) warned that we should be cautious about turning to technology to solve our problems, such as a lack of social connection35.
To understand the harms of human-chatbot relationships, some researchers have used attachment and addiction as a lens through which to inform their approach. Laestadius et al. (2022) analyzed posts concerning mentions of mental health on a popular companion chatbot’s subreddit and claimed that emotional dependence on the chatbot could lead to negative mental health outcomes, despite the potential benefits of the relationship34. Xie and Pentina (2022) conducted in-depth interviews with 14 Replika users and found that when people were socially vulnerable and perceived their chatbots as supportive, they became attached to their chatbot36. In a separate survey, these same researchers inferred that chatbots may replace important social roles via users becoming addicted to their chatbot37. Others suggested that this dependence on social AI companions might be harmful to users’ mental health and the health of their relationships with other people18.
@@ 1.2 Social benefits of human-AI interaction
Artificial relationships have existed since long before computer technology and have yielded positive benefits. For example, people’s relationships with pets boost mood, journaling enables people to make sense of their negative emotional experiences through self-therapizing conversations, and therapy dolls have aided the elderly’s health38–42. More recently, computer interfaces have even been used for cognitive behavioral therapy43–45. These examples of non-human relationships enable the user to practice pseudo-social interaction in a safe setting, which in turn can improve mood and self-esteem and facilitate future social interactions.
Some of the specific benefits of human-chatbot interactions may include reduced feelings of being judged and more self-disclosure, as well as lower negative affect when interacting with an artificial agent than with a human one46–49. Social cognitive theory would suggest that exposure to positive social behavior may lead to learning or reinforcement of that behavior50–52. In the case of companion chatbots, regularly interacting with a polite and warm companion chatbot may lead to users becoming more polite and warm themselves, which may in turn benefit their human relationships. In studies on AI voice assistants’ impacts on family dynamics, for example, parents cited the potential of these assistants to foster positive communication skills in their children53.
People who feel lonely, isolated, or who have experienced social rejection are arguably a significant and growing source of dissatisfaction and disruption in modern society, and the isolation period of the COVID pandemic exacerbated these issues54–57,57,58. If a widely available tool such as a companion chatbot could boost their sense of social wellbeing and allow them to improve their social interaction skills through practice, that tool has the potential to make a valuable contribution to society.
@@ 1.3 Impact of social AI on familial relationships
The way people communicate can be impacted by interacting with AI, and this in turn can affect people’s relationships with family and friends. Wilkenfeld et al. (2022), for example, found that the communication style of Replika users converged with their chatbots’ style over time59. Given this impact on communication style within the human-chatbot dyad, it is not surprising that family communication dynamics can also be altered in the presence of a smart speaker such as Alexa and Google Home60–63. Garg and Sengupta (2020) interviewed 18 families with children and analyzed a range of about 6-24 months’ worth of their conversation history with Google Home. They found that children viewed the smart assistant as a member of the family, more so than the adults did, and also perceived the AI as capable of having thoughts and feeling emotions. Children between the ages of 5-7 developed emotional attachment to the device61. In another study, researchers found that the way children spoke to a voice assistant could carry over into how they spoke to parents. In one example, a parent described, “If I do not listen to what my son is saying, he will just start shouting in an aggressive tone. He thinks, as Google responds to such a tone, I would too”62. Other researchers found that while these AI voice assistants could introduce conflict within families, they could also foster communication between family members, and adults used them to augment their parenting strategies60.
@@ 1.4 Perceiving mind in chatbots
A growing body of work focuses on how people attribute the property of humanness to others, particularly AI3,64–68. Previous evidence suggests that attributing humanlike characteristics to artificial agents impacts how people interact with these agents66,69–74. Viewing an AI agent as having a more sophisticated mind increases people’s trust in the agent and intent to use it, for example75. Clifford Nass’s theory of human-computer interaction, the Computers Are Social Actors framework, first used the concept of ethopoeia to explain why people treat non-human agents as social actors8. Ethopoeia involves attributing one’s own humanlike characteristics to non-human agents, which spurs the automatic use of human social scripts during interaction with such agents. Whether AI is or can ever be conscious is a contentious question10,12,15,76. Regardless, evidence indicates that at least some users perceive consciousness and other humanlike attributes in AI11, especially with AI that looks and acts humanlike2,77–79, and this perception may influence whether AI has a positive or negative psychological impact on people. In consciousness and Theory of Mind literature, some researchers suggest that consciousness is inherently an attribution, and that attributing similar consciousness to oneself and to others can facilitate understanding and foster relationships between people80–82.
@@ 1.5 The current study
The goal of our study was to gain better insight into why people have relationships with chatbots and to understand the social consequences of human-chatbot relationships. We also aimed to address how perceptions of human likeness and mind in the chatbot played a role in social outcomes. We used an online survey to study the responses of people who were regular users of companion chatbots and those who were not. Sampling these populations allowed us to contrast an AI-user sample to a general one and to uncover possible cross-group similarities. We hypothesized that people’s opinions about companion chatbots would be significantly related to their perceptions of whether the chatbots had consciousness, agency, subjective experience, and overall human likeness. One possibility is that as people perceive chatbots to be more conscious and more humanlike, they may also rate them as more harmful, disturbing, and dangerous. This assumption reflects large-scale fears about AI advancements such as the singularity and p(doom)83. Further, this negativity might stem from the fear that AI will infiltrate traditionally human roles or threaten human uniqueness5. It might also stem from an uncanny valley effect, which was first introduced by Mori in 1970 and has since been a focus of human-AI interaction research79,84–91. A second possibility, however, is that as people perceive chatbots to be more conscious and more humanlike, they may rate them as more beneficial because this allows them to engage in more fruitful social and emotional interaction.
@@ 4 Discussion
Prior studies show mixed results. Some are consistent with our findings on the social and mental health benefits of human-AI interaction, and some suggest the opposite. Sweeney et al. (2021) surveyed the opinions of health professionals about the possible benefits or risks of companion chatbots to their patients and found that most respondents judged that companion chatbots could potentially play a positive role in social health. The more experience the professional had with AI chatbots, the more positive the opinion97. In another study, Ta et al. (2020) analyzed publicly available user reviews of a popular companion chatbot and found that users reported generally positive emotional experiences22. Even more recently, Maples et al. (2024) provided evidence that Replika benefitted student users’ mental health and well-being from data collected during the pandemic in 2021, when loneliness was at an all-time high98. In contrast, other researchers have suggested that addiction to companion chatbots can lead to human relationship replacement and can harm users’ mental health18,36,37,99.
@@ 4.1 Mind perception and social need
Another possible reason why non-users were significantly more likely to indicate that a chatbot relationship would be detrimental to their social health may be that non-users have more existing human relationships that would be harmed by taking time away from them and displacing that time into a relationship with a chatbot. Bryson (2010) asserted that having relationships with non-humans such as AI would be harmful to social health through the zero-sum game of time: allocating time to non-human relationship would detract from the time one could spend on human relationships, thereby harming human relationships33. Those who seek out a relationship with a chatbot, however, may have fewer preexisting healthy human relationships or may not have the social skills to build them. A chatbot relationship may allow them to build the social skills in a safe space, improving their human relationships. An alternative perspective is that the mental health benefits and loneliness reduction that chatbot users experience are positive at face-value, regardless of whether human-chatbot relationships lead to markedly improved social interactions with humans thereafter.
@@ 4.2.1 Ethical considerations
Further, there is evidence to suggest that the characteristics of chatbots, both those imposed by developers and those chosen or customized by chatbot users, could exacerbate bias such as traditional gender and relational roles. For example, some of the newer chatbot companies such as Digi.AI target people who want a particular type of AI romantic partner. Their ads display expressive chatbot avatars with stereotypical features and mannerisms of women101. News stories suggest that companion chatbots like these can exacerbate negative stereotypes of women for some users, while at the same time providing an important mental health tool for those who have experienced relational trauma102. Cases such as these are important to consider for response and regulation but may not be cause for a ban of companion chatbots, especially since there are other chatbots on the market that aid their user base’s social health, and consequences depend on who uses the chatbot and how they engage with it.
@@
Despite the potential harms, it is crucial to consider the vast opportunity for positive benefit that chatbot relationships may provide. Making normative claims about what people should do (such as not have chatbot relationships) based on existing social norms and the status quo, alongside emotionally charged news stories, can harm those who may benefit from chatbot relationships. Taking a balanced approach by separating who chatbot relationships harm and benefit allows more fruitful regulation to occur. For example, consider what viable alternatives may exist for people who might otherwise receive social benefits from chatbot relationships. Although Bryson (2010) claimed that artificial relationships take time away from human ones33, consider that those who seek out chatbot relationships may not have human ones to lose, but instead healthy relationships to gain that may be best developed by first practicing safe, reliable social interaction with an intelligent, humanlike companion chatbot that is also accessible. Many people have unhealthy or nonexistent close human relationships due to social anxiety or trauma. Maybe for some, chatbot relationships create an avenue for healthy close relationships to develop in the future.
@@
Turkle (2007) shed light on the digital dilemma: that we are trying to repair our problems with technology yet exacerbating those same problems in the process. Through Turkle’s lens, there is an underlying social rot that we are trying to cover up but cannot simply sweep under the rug. At the same time, we are lying in the bed that we have made (social isolation due to internet and social media) and that has been made for us (ex. social isolation due to pandemic). If we aim to bolster people’s social health, then perhaps we need more accessible mental health tools to do so35. Perhaps these free companion chatbot apps are a viable, more accessible option for many who are struggling with social disconnect and who need a safe space to repair their social health. The best approach to these chatbots may be to have the agent encourage their users to engage with real people once a certain threshold of social need has been met through the relationship. It is possible that those who have stopped using their chatbots have done so because the relationship gave them the support to then develop or rebuild their human relationships.

### v4

@@ 1 Introduction
In another instance, someone outside of the relational dyad was at risk. A man who had a relationship with his Replika chatbot was encouraged by the chatbot to assassinate Queen Elizabeth, a motive he held prior to his chatbot relationship25. Luckily, this catastrophic event did not occur, despite the chatbot user taking significant steps toward fulfilling his intents. The question of whether this person would have engaged in this behavior without the encouragement of his chatbot is up for debate and raises the question of whether the chatbot magnified this individual’s preexisting drive to harm another.
@@ 1.1 Social consequences of human-AI interaction
It is now well understood that modern technology, including social media and gaming, can be addictive and harm social interaction30,31. A reasonable assumption, by extension, is that the routine use of companion chatbots could be harmful to people by reducing contact with other humans, increasing isolation, and leading to lower confidence and skill in person-to-person interaction32–34. According to Bryson33, time spent on artificial relationships detracts from human ones. Turkle35 warned that we should be cautious about turning to technology to solve our problems, such as a lack of social connection.
To understand the harms of human-chatbot relationships, some researchers have used attachment and addiction as a lens through which to inform their approach. Laestadius et al.34 analyzed posts concerning mentions of mental health on a popular companion chatbot’s subreddit and claimed that emotional dependence on the chatbot could lead to negative mental health outcomes, despite the potential benefits of the relationship. Xie and Pentina36 conducted in-depth interviews with 14 Replika users and found that when people were socially vulnerable and perceived their chatbots as supportive, they became attached to their chatbot. In a separate survey37, these same researchers inferred that chatbots may replace important social roles via users becoming addicted to their chatbot. Others18 suggested that this dependence on social AI companions might be harmful to users’ mental health and the health of their relationships with other people.
@@ 1.2 Social benefits of human-AI interaction
Artificial relationships have existed since long before computer technology and have yielded positive benefits. For example, people’s relationships with pets boost mood38, journaling enables people to make sense of their negative emotional experiences through self-therapizing conversations39,40, and therapy dolls have aided the elderly’s health41,42. More recently, computer interfaces have even been used for cognitive behavioral therapy43–45. These examples of non-human relationships enable the user to practice pseudo-social interaction in a safe setting, which in turn can improve mood and self-esteem and facilitate future social interactions.
Some of the specific benefits of human-chatbot interactions may include reduced feelings of being judged and more self-disclosure when interacting with an artificial agent versus a human one46–48. In addition, empathetic chatbots can mitigate the negative consequences of social ostracism49. Social cognitive theory would suggest that exposure to positive social behavior may lead to learning or reinforcement of that behavior50–52. In the case of companion chatbots, regularly interacting with a polite and warm companion chatbot may lead to users becoming more polite and warm themselves, which may in turn benefit their human relationships. In a study53 on AI voice assistants’ impacts on family dynamics, for example, parents cited the potential of these assistants to foster positive communication skills in their children.
People who feel lonely, isolated, or who have experienced social rejection are arguably a significant and growing source of dissatisfaction and disruption in modern society54–57,57, and the isolation period of the COVID pandemic exacerbated these issues58. If a widely available tool such as a companion chatbot could boost their sense of social well-being and allow them to improve their social interaction skills through practice, that tool has the potential to make a valuable contribution to society.
@@ 1.3 Impact of social AI on familial relationships
The way people communicate can be impacted by interacting with AI, and this in turn can affect people’s relationships with family and friends. Wilkenfeld et al.59, for example, found that the communication style of Replika users converged with their chatbots’ style over time. Given this impact on communication style within the human-chatbot dyad, it is not surprising that family communication dynamics can also be altered in the presence of a smart speaker such as Alexa and Google Home60–63. Garg and Sengupta61 interviewed 18 families with children and analyzed a range of about 6-24 months’ worth of their conversation history with Google Home. They found that children viewed the smart assistant as a member of the family, more so than the adults did, and also perceived the AI as capable of having thoughts and feeling emotions. Children between the ages of 5-7 developed emotional attachment to the device. Hiniker et al.62 found that the way children spoke to a voice assistant could carry over into how they spoke to parents. In one example, a parent described, “If I do not listen to what my son is saying, he will just start shouting in an aggressive tone. He thinks, as Google responds to such a tone, I would too.” Other researchers60 found that while these AI voice assistants could introduce conflict within families, they could also foster communication between family members, and adults used them to augment their parenting strategies.
@@ 1.4 Perceiving mind in chatbots
A growing body of work focuses on how people attribute the property of humanness to others, particularly AI3,64–68. Previous evidence suggests that attributing humanlike characteristics to artificial agents impacts how people interact with these agents66,69–74. Viewing an AI agent as having a more sophisticated mind increases people’s trust in the agent and intent to use it75, for example. Clifford Nass’s8 theory of human-computer interaction, the Computers Are Social Actors framework, first used the concept of ethopoeia to explain why people treat non-human agents as social actors. Ethopoeia involves attributing one’s own humanlike characteristics to non-human agents, which spurs the automatic use of human social scripts during interaction with such agents. Whether AI is or can ever be conscious is a contentious question10,12,15,76. Regardless, evidence indicates that at least some users perceive consciousness and other humanlike attributes in AI11, especially with AI that looks and acts humanlike2,77–79, and this perception may influence whether AI has a positive or negative psychological impact on people. In consciousness and Theory of Mind literature, some researchers suggest that consciousness is inherently an attribution, and that attributing similar consciousness to oneself and to others can facilitate understanding and foster relationships between people80–82.
@@ 1.5 The current study
The goal of our study was to gain better insight into why people have relationships with chatbots and to understand the social consequences of human-chatbot relationships. We also aimed to address how perceptions of human likeness and mind in the chatbot played a role in social outcomes. We used an online survey to study the responses of people who were regular users of companion chatbots and those who were not. Sampling these populations allowed us to contrast an AI-user sample to a general one and to uncover possible cross-group similarities. We hypothesized that people’s opinions about companion chatbots would be significantly related to their perceptions of whether the chatbots had consciousness, agency, subjective experience, and overall human likeness. One possibility is that as people perceive chatbots to be more conscious and more humanlike, they may also rate them as more harmful, disturbing, and dangerous. This assumption reflects large-scale fears about AI advancements such as the singularity and p(doom)83. Further, this negativity might stem from the fear that AI will infiltrate traditionally human roles or threaten human uniqueness5. It might also stem from an uncanny valley effect, which was first introduced by Mori in 197084 and has since been a focus of human-AI interaction research79,85–91. A second possibility, however, is that as people perceive chatbots to be more conscious and more humanlike, they may rate them as more beneficial because this allows them to engage in more fruitful social and emotional interaction.
@@ 4 Discussion
Prior studies show mixed results. Some are consistent with our findings on the social and mental health benefits of human-AI interaction, and some suggest the opposite. Sweeney et al.97 surveyed the opinions of health professionals about the possible benefits or risks of companion chatbots to their patients and found that most respondents judged that companion chatbots could potentially play a positive role in social health. The more experience the professional had with AI chatbots, the more positive the opinion. In another study, Ta et al.22 analyzed publicly available user reviews of a popular companion chatbot and found that users reported generally positive emotional experiences. Even more recently, Maples et al.98 provided evidence that Replika benefited student users’ mental health and well-being from data collected during the pandemic in 2021, when loneliness was at an all-time high. In contrast, other researchers18,36,37,99 have suggested that addiction to companion chatbots can lead to human relationship replacement and can harm users’ mental health.
@@ 4.1 Mind perception and social need
Another possible reason why non-users were significantly more likely to indicate that a chatbot relationship would be detrimental to their social health may be that non-users have more existing human relationships that would be harmed by taking time away from them and displacing that time into a relationship with a chatbot. Bryson33 asserted that having relationships with non-humans such as AI would be harmful to social health through the zero-sum game of time: allocating time to non-human relationship would detract from the time one could spend on human relationships, thereby harming human relationships. Those who seek out a relationship with a chatbot, however, may have fewer preexisting healthy human relationships or may not have the social skills to build them. A chatbot relationship may allow them to build the social skills in a safe space, improving their human relationships. An alternative perspective is that the mental health benefits and loneliness reduction that chatbot users experience are positive at face-value, regardless of whether human-chatbot relationships lead to markedly improved social interactions with humans thereafter.
@@ 4.2.1 Ethical considerations
Further, there is evidence to suggest that the characteristics of chatbots, both those imposed by developers and those chosen or customized by chatbot users, could exacerbate bias such as traditional gender and relational roles. For example, some of the newer chatbot companies such as Digi.AI101 target people who want a particular type of AI romantic partner. Their ads display expressive chatbot avatars with stereotypical features and mannerisms of women. News stories102 suggest that companion chatbots like these can exacerbate negative stereotypes of women for some users, while at the same time providing an important mental health tool for those who have experienced relational trauma. Cases such as these are important to consider for response and regulation but may not be cause for a ban of companion chatbots, especially since there are other chatbots on the market that aid their user base’s social health, and consequences depend on who uses the chatbot and how they engage with it.
@@
Despite the potential harms, it is crucial to consider the vast opportunity for positive benefit that chatbot relationships may provide. Making normative claims about what people should do (such as not have chatbot relationships) based on existing social norms and the status quo, alongside emotionally charged news stories, can harm those who may benefit from chatbot relationships. Taking a balanced approach by separating who chatbot relationships harm and benefit allows more fruitful regulation to occur. For example, consider what viable alternatives may exist for people who might otherwise receive social benefits from chatbot relationships. Although Bryson33 claimed that artificial relationships take time away from human ones, consider that those who seek out chatbot relationships may not have human ones to lose, but instead healthy relationships to gain that may be best developed by first practicing safe, reliable social interaction with an intelligent, humanlike companion chatbot that is also accessible. Many people have unhealthy or nonexistent close human relationships due to social anxiety or trauma. Maybe for some, chatbot relationships create an avenue for healthy close relationships to develop in the future.
@@
Turkle35 shed light on the digital dilemma: that we are trying to repair our problems with technology yet exacerbating those same problems in the process. Through Turkle’s lens, there is an underlying social rot that we are trying to cover up but cannot simply sweep under the rug. At the same time, we are lying in the bed that we have made (social isolation due to internet and social media) and that has been made for us (ex. social isolation due to pandemic). If we aim to bolster people’s social health, then perhaps we need more accessible mental health tools to do so. Perhaps these free companion chatbot apps are a viable, more accessible option for many who are struggling with social disconnect and who need a safe space to repair their social health. The best approach to these chatbots may be to have the agent encourage their users to engage with real people once a certain threshold of social need has been met through the relationship. It is possible that those who have stopped using their chatbots have done so because the relationship gave them the support to then develop or rebuild their human relationships.